A sealed container was filled with 0.300 mol H2(g), 0.400 mol I2(g), and 0.200 mol HI(g) at 870 K and total pressure 1.00 bar. Calculate the amounts of the components in the mixture at equilibrium given that K = 870 for the reaction H2(g) + I2(g) equilibrium reaction arrow 2 HI(g).
1
Expert's answer
2019-12-05T08:19:31-0500

Solution.

"p \times V = n \times R \times T"

"V = \frac{n \times R \times T}{p}"

V = 65.06 L

"C(I2) = 0.0062 \ M"

"C(H2) = 0.0046 \ M"

"C(HI) = 0.0031 \ M"

"Kp = Kc \times (R \times T)^{(- \Delta \nu)}"

"For \ this \ reaction \ \Delta \nu = 0 and \ Kp = Kc"

"Kp = \frac{[HI]^2eq}{]H2]eq \times [I2]eq}"

"870 = \frac{(0.0031+x)^2}{(0.0046-x) \times (0.0062-x)}"

x = 0.0045

"[HI] = (0.0031+2\times 0.0045) \times 65.06 = 0.79 \ mol"

"[I2] = (0.0062-0.0045) \times 65.06 = 0.11 \ mol"

"[H2] = (0.0046-0.0045) \times 65.06 = 0.0065 \ mol"

Answer:

"[HI] = 0.79 \ mol"

"[I2] = 0.11 \ mol"

"[H2] = 0.0065 \ mol"